if a number is increased by 30, it is less than its square by 12. Find the number
Answer by stanbon(75887)

if a number is increased by 30, it is less than its square by 12. Find the number
---------------
x + 30 = x^2-12
-------
x^2 - x -42 = 0
========
Factor:
(x-7)(x+6) = 0
---
x = 7 or x = -6
===================
Cheers,
Stan H.
===================